# Probing nuclear liquid-gas phase transition with isospin correlations

### v1

@@ Abstract
We investigate the fluctuations of the net-baryon number near the critical point of the liquid-gas phase transition. We use the parity doublet model in the mean-field approximation fixed to the zero-temperature properties of nuclear matter to account for critical behavior. We explicitly calculate the fluctuations of the net-proton and net-neutron numbers as well as their correlations in the isospin-symmetric matter. We focus on the qualitative properties and systematics of the first- to fourth-order susceptibilities and their ratios. We demonstrate that the fluctuations of net-proton number do not reflect the total net-baryon number fluctuations in the vicinity of the nuclear liquid-gas phase transition. We point to the importance of the non-trivial correlations between protons and neutrons.
@@ I Introduction
Locating a putative QCD critical point is part of the experimental effort in large-scale heavy-ion collision experiments. It was suggested that the phase structure of QCD can be investigated through fluctuations and correlations of conserved charges. They are known to be promising theoretical observables in search of critical behavior at the QCD phase boundary [18, 19, 20, 21]. In particular, fluctuations of conserved charges have been proposed to probe the QCD critical point, as well as the remnants of the $O(4)$ criticality at vanishing and finite net-baryon densities [21, 22, 23, 24, 25]. Non-monotonic behavior is expected for various ratios of the cumulants of the net-baryon number. The results of the Beam Energy Scan at the Relativistic Heavy Ion Collider, which covered $\sqrt{s_{\rm NN}}=3-200~{}$GeV, have shown indications of a nonmonotonic behavior in cumulants of the net-proton multiplicity distributions [26, 27, 28, 29, 30, 31]. However, no definitive consensus has been reached regarding the existence and location of the QCD critical point. More data and higher statistics at low collision energies are needed to draw firm conclusions.
@@
It has been argued that at large baryon chemical potentials, the chemical freeze-out curve should approach the critical point of the nuclear liquid-gas phase transition [32]. The HADES Collaboration at GSI will probe fluctuations at low energies $0.2-1.0A~{}$GeV, with the expectation of exploring the nuclear liquid-gas phase transition region [33]. The properties of nuclear matter and the liquid-gas phase transition have been studied in the literature (see, e.g., [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]).
@@
Due to experimental limitations, only charged particles created in a heavy-ion collision have a fair chance of being measured by the detector. The net-proton number fluctuations are assumed to reliably reflect the net-baryon number fluctuations because proton is the lightest and most abundant baryon created in heavy-ion collisions. This simplification works at high to intermediate energies (see, e.g., [48]). However, at low energies, where the proton distribution is dominated by interactions between hadrons [30], the net-proton number fluctuations could cease to be a good proxy for the net-baryon number fluctuations. Qualitative differences in net-baryon number density fluctuations and individual baryonic chiral partners of opposite parity were examined near the nuclear liquid-gas and chiral phase transitions [45, 44, 49]. However, the relations and differences between net-baryon and net-proton number fluctuations have not yet been explored enough in theoretical models near the liquid-gas phase transition.
@@
In this work, we apply the parity doublet model [50, 51, 52] to calculate the generalized susceptibilities of the net-baryon number density. Specifically, we focus on the fluctuations of protons and neutrons, as well as correlations among them. We calculate susceptibilities up to the fourth order and examine their qualitative behavior near the nuclear liquid-gas phase transition.
II Fluctautions near the nuclear liquid-gas phase transition
@@
Figure 1: Net-baryon number susceptibility at $T=30~{}\rm MeV$ in the vicinity of the liquid-gas phase transition. The gray, dotted vertical line marks the crossover transition obtained from the maximum of $\hat{\chi}_{2}^{B}$.
@@
where $\Omega_{x}$’s are the kinetic parts of the thermodynamic potential for positive-parity nucleons, protons ($p$) and neutrons ($n)$ identified with $N(939)$, and their negative-parity chiral partners, $p^{\star}$ and $n^{\star}$, identified with $N^{\star}(1535)$ [53]. The potentials $V_{\sigma/\omega}$ are introduced for the scalar ($\sigma$) and vector ($\omega$) mean fields. In-medium profiles of the mean fields are obtained by extremizing the thermodynamic potential in the direction of the mean fields, leading to the gap equations. The set of parameters used in this work and a detailed description of the model can be found in [44].
At low temperatures, the model predicts sequential first-order nuclear liquid-gas and chiral phase transitions with critical points located at $T_{c}=16~{}\rm MeV$, $\mu_{B}=909~{}\rm MeV$ ($n_{B}=0.053~{}{\rm fm^{-3}}=0.33~{}n_{\rm sat}$) and $T_{c}=7~{}\rm MeV$, $\mu_{B}=1526~{}\rm MeV$ ($n_{B}=1.25{\rm~{}fm^{-3}}=7.82~{}n_{0}$), respectively. Here, we pay attention to the fluctuations near the critical point of the liquid-gas phase transition. Thus, we restrict our considerations to low temperatures and small densities. The vacuum mass of the negative-parity nucleons is equal to 1.51.51.5~{}1.5GeV [53]. Effectively, the negative-parity nucleons are thermally suppressed near the liquid-gas phase transition, and the system is predominantly composed of positive-parity nucleons [44, 45]. Therefore, in this region, the thermodynamic potential can be reliably restricted to
@@
Figure 3: Third-order (left panel) and fourth-order (right panel) susceptibilities of the net-baryon number at $T=30~{}\rm MeV$ in the vicinity of the liquid-gas phase transition. The gray, dotted vertical line marks the crossover transition obtained from the maximum of $\hat{\chi}_{2}^{B}$.
@@
Fig. 1 shows the susceptibility of the net-baryon number density, $\hat{\chi}_{2}^{B}$, at $T=30~{}\rm MeV$ in the vicinity of the liquid-gas phase transition. It features a clear maximum as it passes through the crossover transition and decreases after it. Initially, at small baryon chemical potential, the susceptibility $\hat{\chi}_{2}^{p}$ also increases and features a maximum in the vicinity of the crossover transition, followed by a shallow minimum, after which it continues to increase. Notably, as the crossover transition is approached, $\hat{\chi}_{2}^{p}$ starts to deviate from $\hat{\chi}_{2}^{B}$ and has an opposite trend after the transition. This already indicates that the correlator $\hat{\chi}_{2}^{pn}$ is non-vanishing and has a non-trivial structure [cf. Eq. (6)]. In fact, we find that at small chemical potentials, $\hat{\chi}_{2}^{pn}\simeq 0$ and $\hat{\chi}_{2}^{B}\simeq 2\hat{\chi}_{2}^{p}$. However, as the crossover transition is approached, the correlator increases, and a peak is observed in the vicinity of the transition, and it continues to decrease, eventually becoming negative. This causes an enhancement of $\hat{\chi}_{2}^{p}$ compared to $\hat{\chi}_{2}^{B}$ after the transition. It should be noted that experimentally measurable fluctuations of the proton number are identified as proxies of the baryon number fluctuations. From Eq. (7), it is clear that such an assumption requires a vanishing $\hat{\chi}_{2}^{pn}$ correlator. Our results indicate that this is not the case, especially in the vicinity of a CP. In fact, susceptibilities such as $\hat{\chi}_{2}^{ud}\propto\hat{\chi}_{2}^{pn}$ are not vanishing at finite temperature in LQCD simulations [55] and model calculations [56, 57]. Here, finite proton-neutron correlations appear due to the presence of interactions mediated by the mean fields [cf. (5)].
In Fig. 2, we show the behavior of the correlator $\hat{\chi}_{2}^{pn}$ in the $\mu_{B}-T$ plane. At any temperature shown, $\hat{\chi}_{2}^{pn}\rightarrow 0$ as $\mu_{B}$ approaches zero, indicating that the system behaves as a free gas of nucleons. In the vicinity of a CP, the correlator increases substantially, develops a maximum, and turns negative at larger chemical potentials. Due to the nontrivial structure of the $\hat{\chi}_{2}^{pn}$ correlator near the CP, a common identification of the net-proton number fluctuations as half of the net-baryon number fluctuations is violated. This result is obtained in the isospin-symmetric case, i.e., $n_{B}=2~{}n_{p/n}$, with complete randomization of the isospin space.
@@
respectively. The above definitions are direct generalizations of Eqs. (7) and (8). In our case, $\hat{\chi}_{k}^{p}=\hat{\chi}_{k}^{n}$ due to isospin symmetry. We note that each term in Eq. (10) could be studied in principle individually, revealing more complex structures of the correlator. We plan to elaborate on this elsewhere.
Fig. 3 shows the third and fourth-order susceptibilities at $T=30~{}$MeV. Both susceptibilities of the net-baryon number density ($\hat{\chi}_{3}^{B}$ and $\hat{\chi}_{4}^{B}$) exhibit distinctive behavior and show strong sensitivity to dynamical effects related to the liquid-gas phase transition. Interestingly, we observe that susceptibilities $\hat{\chi}_{3,4}^{p}$ do not qualitatively reflect $\hat{\chi}_{3,4}^{B}$. The third-order susceptibility $\hat{\chi}_{3}^{p}$ is positive and shows a maximum in the vicinity of the crossover transition. The fourth-order susceptibility $\hat{\chi}_{4}^{p}$ features a peak-dip structure. However, it does not turn positive after the crossover transition, compared to $\hat{\chi}_{4}^{B}$. In contrast to this, the correlators $\hat{\chi}_{3}^{pn}$ and $\hat{\chi}_{4}^{pn}$ have structures that qualitatively resemble those of $\hat{\chi}_{3}^{B}$ and $\hat{\chi}_{4}^{B}$, respectively.
@@
III Summary

### v2

@@ Abstract
We investigate the fluctuations of the net-baryon number near the critical point of the liquid-gas phase transition. We use the parity doublet model in the mean-field approximation fixed to the zero-temperature properties of nuclear matter to account for critical behavior. We explicitly calculate the fluctuations of the net-proton and net-neutron numbers as well as their correlations in the isospin-symmetric matter. We focus on the qualitative properties and systematics of the first- to fourth-order susceptibilities and their ratios. We demonstrate that the fluctuations of net-proton number do not reflect the total net-baryon number fluctuations in the vicinity of the nuclear liquid-gas phase transition. We also study the behavior of the baryon and proton number factorial cumulants. We highlight the importance of the non-trivial correlations between protons and neutrons.
@@ I Introduction
Locating a putative QCD critical point is part of the experimental effort in large-scale heavy-ion collision experiments. It was suggested that the phase structure of QCD can be investigated through fluctuations and correlations of conserved charges. They are known to be promising theoretical observables in search of critical behavior at the QCD phase boundary [18, 19, 20, 21]. In particular, fluctuations of conserved charges have been proposed to probe the QCD critical point, as well as the remnants of the $O(4)$ criticality at vanishing and finite net-baryon densities [21, 22, 23, 24, 25]. Non-monotonic behavior is expected for various ratios of the cumulants of the net-baryon number. The results of the Beam Energy Scan at the Relativistic Heavy Ion Collider, which covered $\sqrt{s_{\rm NN}}=3-200$~GeV, have shown indications of a non-monotonic behavior in cumulants of the net-proton multiplicity distributions [26, 27, 28, 29, 30, 31]. However, no definitive consensus has been reached regarding the existence and location of the QCD critical point. More data and higher statistics at low collision energies are needed to draw firm conclusions.
@@
It has been argued that at large baryon chemical potentials, the chemical freeze-out curve should approach the critical point of the nuclear liquid-gas phase transition [32]. The HADES Collaboration at GSI will probe fluctuations at low energies 0.21.0A0.2-1.0A~GeV, with the expectation of exploring the nuclear liquid-gas phase transition region [33]. The properties of nuclear matter and the liquid-gas phase transition have been studied in the literature (see, e.g., [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]).
@@
Due to experimental limitations, only charged particles created in a heavy-ion collision have a fair chance of being measured by the detector. The net-proton number fluctuations are assumed to reliably reflect the net-baryon number fluctuations because proton is the lightest and most abundant baryon created in heavy-ion collisions (HICs). This simplification works for susceptibilities calculated in an ideal gas formulated in a grand-canonical ensemble [48]. However, off-diagonal components of susceptibilities can be induced by the presence of interactions [44]. In HICs, additional effects, such as exact conservation of the baryon number [49, 24, 50], can cause differences between the net-proton and net-baryon cumulants [51]. The baryon cumulants can be reconstructed from the proton cumulants in the case of full randomization of isospin under the assumption that the quantum correlations arising from Fermi statistics are negligible [52]. However, this is not the case at low energies, where the proton distribution is dominated by interactions between hadrons [30], the net-proton number fluctuations could cease to be a good proxy for the net-baryon number fluctuations. Qualitative differences in the net-baryon number density fluctuations and individual baryonic chiral partners of opposite parity were examined near the nuclear liquid-gas and chiral phase transitions [45, 44, 53]. However, the relations and differences between net-baryon and net-proton number fluctuations have not yet been explored enough in theoretical models near the liquid-gas phase transition.
@@
In this work, we apply the parity doublet model [54, 55, 56] to calculate the generalized susceptibilities of the net-baryon number density and the factorial cumulants of the baryon number. Specifically, we focus on the fluctuations of protons and neutrons, as well as the correlations among them. We calculate susceptibilities and factorial cumulants up to the fourth order and examine their qualitative behavior near the nuclear liquid-gas phase transition.
II Fluctuations near the nuclear liquid-gas phase transition
@@
Figure 1: Net-baryon number susceptibility at $T=30~\rm MeV$ in the vicinity of the liquid-gas phase transition. The grey, dotted vertical line marks the crossover transition obtained from the maximum of $\hat{\chi}_{2}^{B}$.
@@
where $\Omega_{x}$’s are the kinetic parts of the thermodynamic potential for positive-parity nucleons, protons ($p$) and neutrons ($n)$ identified with $N(939)$, and their negative-parity chiral partners, $p^{\star}$ and $n^{\star}$, identified with $N^{\star}(1535)$ [57]. The potentials $V_{\sigma/\omega}$ are introduced for the scalar ($\sigma$) and vector ($\omega$) mean fields. In-medium profiles of the mean fields are obtained by extremizing the thermodynamic potential in the direction of the mean fields, leading to the gap equations. The set of parameters used in this work and a detailed model description can be found in [44].
At low temperatures, the model predicts sequential first-order nuclear liquid-gas and chiral phase transitions with critical points located at $T_{c}=16~\rm MeV$, $\mu_{B}=909~\rm MeV$ ($n_{B}=0.053~{\rm fm^{-3}}=0.33~n_{\rm sat}$) and $T_{c}=7~\rm MeV$, $\mu_{B}=1526~\rm MeV$ ($n_{B}=1.25{\rm~fm^{-3}}=7.82~n_{0}$), respectively. Here, we pay attention to the fluctuations near the critical point of the liquid-gas phase transition. Thus, we restrict our considerations to low temperatures and small densities. The vacuum mass of the negative-parity nucleons equals 1.51.5~GeV [57]. Effectively, the negative-parity nucleons are thermally suppressed near the liquid-gas phase transition, and the system is predominantly composed of positive-parity nucleons [44, 45]. Therefore, in this region, the thermodynamic potential can be reliably restricted to
@@
Figure 3: Third-order (left panel) and fourth-order (right panel) susceptibilities of the net-baryon number at $T=30~\rm MeV$ in the vicinity of the liquid-gas phase transition. The grey, dotted vertical line marks the crossover transition obtained from the maximum of $\hat{\chi}_{2}^{B}$.
@@
Fig. 1 shows the susceptibility of the net-baryon number density, $\hat{\chi}_{2}^{B}$, at $T=30~\rm MeV$ in the vicinity of the liquid-gas phase transition. It features a clear maximum as it passes through the crossover transition and decreases afterward. Initially, at small baryon chemical potential, the susceptibility $\hat{\chi}_{2}^{p}$ also increases and features a maximum in the vicinity of the crossover transition, followed by a shallow minimum, after which it continues to grow. Notably, as the crossover transition is approached, $\hat{\chi}_{2}^{p}$ starts to deviate from $\hat{\chi}_{2}^{B}$ and has an opposite trend after the transition. This already indicates that the correlator $\hat{\chi}_{2}^{pn}$ is non-vanishing and has a non-trivial structure [cf. Eq. (6)]. We find that at small chemical potentials, $\hat{\chi}_{2}^{pn}\simeq 0$ and $\hat{\chi}_{2}^{B}\simeq 2\hat{\chi}_{2}^{p}$. However, as the crossover transition is approached, the correlator increases, a peak is observed in the vicinity of the transition, and it continues to decrease, eventually becoming negative. This causes an enhancement of $\hat{\chi}_{2}^{p}$ compared to $\hat{\chi}_{2}^{B}$ after the transition. It should be noted that experimentally measurable fluctuations of the proton number are identified as proxies of the baryon number fluctuations. From Eq. (7), it is clear that such an assumption requires a vanishing $\hat{\chi}_{2}^{pn}$ correlator. Our results indicate that this is not the case, especially in the vicinity of a CP. In fact, susceptibilities such as the up-down quark correlator $\hat{\chi}_{2}^{ud}$ are not vanishing at finite temperature in LQCD simulations [59] and model calculations [60, 61].
Using $\mu_{p}=2\mu_{u}+\mu_{d}$ and $\mu_{n}=\mu_{u}+2\mu_{d}$, the correlator $\hat{\chi}_{2}^{ud}$ can be expressed as
$\hat{\chi}_{2}^{ud}=4\hat{\chi}_{2}^{p}+5\hat{\chi}_{2}^{pn}\textrm{,}$ (9)
where we have assumed that $\hat{\chi}_{2}^{p}=\hat{\chi}_{2}^{n}$. Even in the absence of correlations, $\hat{\chi}_{2}^{ud}\sim\hat{\chi}_{2}^{p}$ is non-vanishing due to the quark substructure of nucleons [62]. Any qualitative difference between $\hat{\chi}_{2}^{ud}$ and $\hat{\chi}_{2}^{p}$ amounts to the non-trivial proton-neutron correlations. Here, they appear due to the presence of interactions mediated by the mean fields [cf. (5)]. We point out that in a more general case mesons contribute to $\hat{\chi}_{2}^{ud}$ as well with negative sign [59] at finite temperature.
In Fig. 2, we show the behavior of the correlator $\hat{\chi}_{2}^{pn}$ in the $\mu_{B}-T$ plane. At any temperature shown, $\hat{\chi}_{2}^{pn}\rightarrow 0$ as $\mu_{B}$ approaches zero, indicating that the system behaves as a free gas of nucleons. In the vicinity of a CP, the correlator increases substantially, develops a maximum, and turns negative at larger chemical potentials. Due to the non-trivial structure of the $\hat{\chi}_{2}^{pn}$ correlator near the CP, a common identification of the net-proton number fluctuations as half of the net-baryon number fluctuations is violated. This result is obtained in the isospin-symmetric case, i.e., $n_{B}=2~n_{p/n}$, with complete randomization of the isospin space.
@@
respectively. The above definitions are direct generalizations of Eqs. (7) and (8), and
$\hat{\chi}_{k}^{B}=\hat{\chi}_{k}^{p}+\hat{\chi}_{k}^{n}+\hat{\chi}_{k}^{pn}\textrm{.}$ (12)
In our case, $\hat{\chi}_{k}^{p}=\hat{\chi}_{k}^{n}$ due to isospin symmetry. We note that each term in Eq. (11) could be studied in principle individually, revealing more complex structures of the correlator. We plan to elaborate on this elsewhere.
Fig. 3 shows the third and fourth-order susceptibilities at T=30T=30~MeV. Both susceptibilities of the net-baryon number density ($\hat{\chi}_{3}^{B}$ and $\hat{\chi}_{4}^{B}$) exhibit distinctive behavior and show strong sensitivity to dynamical effects related to the liquid-gas phase transition. Interestingly, we observe that susceptibilities $\hat{\chi}_{3,4}^{p}$ do not qualitatively reflect $\hat{\chi}_{3,4}^{B}$. The third-order susceptibility $\hat{\chi}_{3}^{p}$ is positive and shows a maximum in the vicinity of the crossover transition. The fourth-order susceptibility $\hat{\chi}_{4}^{p}$ features a peak-dip structure. However, it does not turn positive after the crossover transition, compared to $\hat{\chi}_{4}^{B}$. In contrast to this, the correlators $\hat{\chi}_{3}^{pn}$ and $\hat{\chi}_{4}^{pn}$ have structures that qualitatively resemble those of $\hat{\chi}_{3}^{B}$ and $\hat{\chi}_{4}^{B}$, respectively.
Figure 5: Second-order factorial cumulants, $\hat{C}^{\alpha}_{2}$, at $T=30~\rm MeV$ in the vicinity of the liquid-gas phase transition. The grey, dotted vertical line marks the crossover transition obtained from the maximum of $\hat{\chi}_{2}^{B}$.
@@
Figure 6: Third-order (left panel) and fourth-order (right panel) factorial cumulants of the baryon and proton numbers at $T=30~\rm MeV$ in the vicinity of the liquid-gas phase transition. The grey, dotted vertical line marks the crossover transition obtained from the maximum of $\hat{\chi}_{2}^{B}$.
In a two-component system of protons and neutrons, the fluctuations of conserved charges, net-baryon ($B$) and net-charge ($Q$) numbers can be identified with specific mixtures of $\hat{\chi}_{2}^{p/n}$. The relations for the net-baryon number susceptibilities are given in Eqs. (7) and (12). For the second-order charge fluctuations and baryon-charge correlator one gets
$\hat{\chi}_{2}^{Q}$ $=-\frac{\mathrm{d}^{2}\hat{\Omega}}{\mathrm{d}\hat{\mu}_{Q}^{2}}\Bigg{|}_{T}=\hat{\chi}_{2}^{p}\textrm{,}$ (14)
$\hat{\chi}_{2}^{BQ}$ $=-\frac{\mathrm{d}^{2}\hat{\Omega}}{\mathrm{d}\hat{\mu}_{B}\mathrm{d}\hat{\mu}_{Q}}\Bigg{|}_{T}=\hat{\chi}_{2}^{p}+\frac{1}{2}\hat{\chi}_{2}^{pn}=\frac{1}{2}\hat{\chi}_{2}^{B}\textrm{,}$ (15)
respectively. We note that the last equality in Eq. (15) is valid due to isospin symmetry. Similarly for the third- and fourth-order susceptibilities
$\hat{\chi}_{3}^{Q}$ $=-\frac{\mathrm{d}^{3}\hat{\Omega}}{\mathrm{d}\hat{\mu}_{Q}^{3}}\Bigg{|}_{T}=\hat{\chi}_{3}^{p}$ (16)
$\hat{\chi}_{4}^{Q}$ $=-\frac{\mathrm{d}^{4}\hat{\Omega}}{\mathrm{d}\hat{\mu}_{Q}^{4}}\Bigg{|}_{T}=\hat{\chi}_{4}^{p}$ (17)
We have verified the above relations also numerically. Various baryon-charge correlators at higher orders are given by a mixture of correlation functions. We elaborate on these in our future work.
In the vicinity of the liquid-gas phase transition, the nuclear medium is predominantly comprised of protons and neutrons. Therefore, one expects that, $\hat{\chi}_{k}^{Q}\simeq\hat{\chi}_{k}^{p}$, as derived in Eqs. (14) and (16). Moreover, the results obtained in our work qualitatively agree with other approaches to nuclear liquid-gas phase transition based on the quantum van der Waals interactions [65]. However, we note that the above does not need to be the case at higher temperature and/or baryon chemical potential, where contributions from the other baryons and their resonances are to be expected.
We note that the nuclear symmetry energy controls the change of the binding energy of nuclear matter as the proton-to-neutron ratio changes. Thus, higher values of the symmetry energy should suppress the individual proton and neutron fluctuations and enhance the correlations among them. However, in the present work, we have not considered the effect of the symmetry energy. Phenomenologically, to fit the physical value for the symmetry energy, the $\rho$ meson coupling to nucleons has to be added to the model Lagrangian. We leave this interesting and important issue for our future work.
III Factorial Cumulants
Baryon number cumulants can be reconstructed from proton cumulants measured experimentally under the assumption of total isospin randomization in the final state of a heavy-ion collision [52]. In this context, useful are factorial cumulants ($\hat{C}_{n}$) as they eliminate self-correlations. Then, the proton number factorial cumulants $\hat{C}_{n}^{p}$ are directly proportional to baryon number factorial cumulants ($\hat{C}^{B}_{n}$),
$\hat{C}_{n}^{B}=2^{n}\hat{C}_{n}^{p}\textrm{.}$ (18)
However, it should be noted that this proportionality is only valid under the assumption that the correlations arising from quantum Fermi statistics are negligible. This is justified at high energies $\sqrt{s_{NN}}\gtrsim 10$~GeV [52], which corresponds to the freeze-out temperature $T\gtrsim 150$~MeV [66]. This is not necessarily the case at low energies, where the proton distribution is dominated by interactions between hadrons [30], and especially in the vicinity of the nuclear liquid-gas phase transition, where Fermi statistics effects are certainly non-negligible.
Factorial cumulants can be expressed as linear combinations of the ordinary susceptibilities [67, 68, 69]. The first four factorial cumulants read:
$\hat{C}^{\alpha}_{1}$ $=\hat{\chi}_{1}^{\alpha}\textrm{,}$ (19)
$\hat{C}^{\alpha}_{2}$ $=\hat{\chi}_{2}^{\alpha}-\hat{\chi}_{1}^{\alpha}\textrm{,}$ (20)
$\hat{C}^{\alpha}_{3}$ $=\hat{\chi}_{3}^{\alpha}-3\hat{\chi}_{2}^{\alpha}+2\hat{\chi}_{1}^{\alpha}\textrm{,}$ (21)
$\hat{C}^{\alpha}_{4}$ $=\hat{\chi}_{4}^{\alpha}-6\hat{\chi}_{3}^{\alpha}+11\hat{\chi}_{2}^{\alpha}-6\hat{\chi}_{1}^{\alpha}\textrm{,}$ (22)
where $\alpha=B,p,n$. Typically, factorial cumulants, $\hat{C}_{n}$, are defined as linear combinations of ordinary cumulants, $\hat{\kappa}_{n}$, which are related to the susceptibilities through $\kappa_{n}=VT^{3}\hat{\chi}_{n}$. Therefore, the definition of factorial cumulants in Eqs. (19)-(22) differ by the factor $VT^{3}$ from usual definition through ordinary cumulants. Nevertheless, our definition does not lead to qualitative differences and the prefactors cancel out when ratios of cumulants are considered. We note that the correlation functions $\hat{C}^{\alpha}_{k}$ are defined for particles and not the net-particle numbers. However, in the vicinity of the liquid-gas phase transition, the contribution of antiparticles is well suppressed and we can stick to the relations above. Nevertheless, a direct relation between the factorial cumulants and susceptibilities for net-particle numbers can also be derived [68, 69]. In the isospin-symmetric system, $\hat{C}_{k}^{p}=\hat{C}_{k}^{n}$.
Figure 7: Ratio $F_{3,1}^{\alpha}$ (top panel) and $F_{4,1}^{\alpha}$ (bottom panel) for factorial cumulants of baryons (left panel) and protons (right panel). The black, solid line marks the first-order liquid-gas phase transition ending with a circle that indicates the critical point. The black, dotted line indicates the crossover transition obtained from the maximum of $\hat{\chi}_{2}^{B}$.
The second-order factorial cumulant for the baryon number at T=30T=30~MeV is depicted in Fig. 5. It develops a peak around the remnant of the liquid-gas phase transition and turns negative, driven by the repulsive interactions. Notably, the factorial cumulant for protons, $\hat{C}_{2}^{p}$, resembles qualitatively the same structure. Although $\hat{C}_{2}^{B}\simeq C_{2}^{p}$, we find that their difference has a non-trivial structure as well. The difference between baryon number factorial cumulants and the factorial cumulants for individual species, which we denote as $\hat{C}_{k}^{pn}$, are given as follows:
$\hat{C}_{1}^{pn}\equiv\hat{C}_{1}^{B}-\sum_{\alpha=p,n}\hat{C}_{1}^{\alpha}$ $=0\textrm{,}$ (23)
$\hat{C}_{2}^{pn}\equiv\hat{C}_{2}^{B}-\sum_{\alpha=p,n}\hat{C}_{2}^{\alpha}$ $=\hat{\chi}_{2}^{pn}\textrm{,}$ (24)
$\hat{C}_{3}^{pn}\equiv\hat{C}_{3}^{B}-\sum_{\alpha=p,n}\hat{C}_{3}^{\alpha}$ $=\hat{\chi}_{3}^{pn}-3\hat{\chi}_{2}^{pn}\textrm{,}$ (25)
$\hat{C}_{4}^{pn}\equiv\hat{C}_{4}^{B}-\sum_{\alpha=p,n}\hat{C}_{4}^{\alpha}$ $=\hat{\chi}_{4}^{pn}-6\hat{\chi}_{3}^{pn}+11\hat{\chi}_{2}^{pn}\textrm{.}$ (26)
From the above equations, one sees that the difference between $\hat{C}_{2}^{B}$ and $\hat{C}_{2}^{p}$ amounts to the proton-neutron correlator, $\hat{\chi}_{2}^{pn}$. Departure of $\hat{C}_{k}^{pn}$ from zero, signals physics beyond the uncorrelated gas of nucleons, which is expected in the vicinity of the liquid-gas phase transition. We note that the relations in Eqs. (23)-(26) do not require additional assumptions and are expected to work for any temperature and chemical potential. In general $\hat{C}_{k}^{pn}\neq 0$ breaks the scaling in Eq. (18). Nevertheless, it is possible to retain it, if $\hat{C}_{k}^{pn}=\left(2^{k}-2\right)\hat{C}_{k}^{p}$. However, we do not find this to be the case in general and $\hat{C}_{k}^{pn}$ is not proportional to $\hat{C}_{k}^{p}$ [cf. Figs. 5 and 6].
In Fig. 6, we show the third- and fourth-order factorial cumulants. The proton number factorial cumulants behave qualitatively similar to the baryon number factorial cumulants. Interestingly, the difference between them show the same qualitative features as $\hat{C}_{k}^{B}$ and $\hat{C}_{k}^{p}$. This is in clear contrast to the susceptibilities [cf. Figs. 1 and 3].
Similarly to the susceptibilities, we define ratios of factorial cumulants,
$F_{n,k}^{\alpha}=\frac{\hat{C}^{\alpha}_{n}}{\hat{C}^{\alpha}_{k}}$ (27)
for $\alpha=B,p,n$. In Fig. 7, we show the ratios of the factorial cumulants $F_{3,1}^{\alpha}$ and $F_{4,1}^{\alpha}$. The structure of factorial-cumulant ratios for protons and baryons remains qualitatively the same for $F_{3,1}^{\alpha}$ and $F_{4,1}^{\alpha}$. Nevertheless, their differences can be shown to be directly proportional to the mixture of correlation functions,
$F_{2,1}^{B}-F_{2,1}^{p}$ $=\frac{\hat{C}_{2}^{pn}}{\hat{C}_{1}^{B}}\textrm{,}$ (28)
$F_{3,1}^{B}-F_{3,1}^{p}$ $=\frac{\hat{C}_{3}^{pn}}{\hat{C}_{1}^{B}}\textrm{,}$ (29)
$F_{4,1}^{B}-F_{4,1}^{p}$ $=\frac{\hat{C}_{4}^{pn}}{\hat{C}_{1}^{B}}\textrm{,}$ (30)
where we have assumed isospin symmetry. The differences between the ratios are directly proportional to the differences derived in Eqs. (23)-(26).
Factorial cumulants eliminate self-correlations that arise from trivial contributions in regular susceptibilities (or cumulants). This allows us to identify the genuine correlations in the system. We find that the presence of interactions, i.e., $\hat{\chi}_{k}^{pn}\neq 0$, modifies the scaling $\hat{C}_{k}^{B}=2^{k}\hat{C}_{k}^{p}$, which is expected since the difference between $\hat{C}_{k}^{B}$and $\hat{C}_{k}^{p}$ is directly proportional to the linear combination of correlations up to $k$-th order. We find that although the baryon and proton cumulants have qualitatively the same structure in the vicinity of the liquid-gas phase transition, their difference is also non-trivial, showing similar features.
IV Summary